# 2]Access e.V., Aachen, Germany

### v1

@@ Abstract
In this study, we develop a novel multi-fidelity deep learning approach that transforms low-fidelity solution maps into high-fidelity ones by incorporating parametric space information into a standard autoencoder architecture. It is shown that, due to the integration of parametric space data, this method requires significantly less training data to achieve effective performance in predicting high-fidelity solution from the low-fidelity one. Consequently, users can conduct their calculations in a lower-dimensional space in a more cost-effective manner, and then upscale the low-resolution solution to a higher-fidelity one without overlooking any significant details. In this study, our focus is on a two-dimensional steady-state heat transfer analysis in highly heterogeneous materials microstructure, where the spatial distribution of heat conductivity coefficients for two distinct materials is condensed from a $101\times 101$ grid to grids of various sizes. Subsequently, the boundary value problem is solved on the coarsest grid using a pre-trained physics-informed neural operator network called as Finite Operator Learning (FOL). Afterward, the calculated low-fidelity result is upscaled to a $101\times 101$ grid using the newly designed enhanced autoencoder. The novelty of the developed enhanced autoencoder lies in the concatenation of heat conductivity maps of different resolutions to the decoder segment in distinct steps. Hence the developed algorithm is named microstructure-embedded autoencoder (MEA). We then compare the outcomes of MEA with the corresponding finite element results, standard U-Net architecture as well as other upscaling approaches such as interpolation functions of varying orders and feedforward neural networks (FFNN). The analysis of the results based on the new approach demonstrates superior performance compared to other approaches in terms of computational cost and error on the test cases. Therefore, as a potential supplement to neural operators networks, our MEA architecture upscales low-fidelity solutions to high-fidelity ones while preserving critical details that are often lost in conventional upscaling methods, especially at sharp interfaces, such as those encountered with interpolation methods.
@@ 1 Introduction
To address the challenges associated with multi-fidelity approaches, we introduced a novel architecture called the Microstructure-Embedded Autoencoder (MEA). In this architecture, the low-fidelity solution map is first obtained by a solver on the reduced parametric space. Afterwards, the resulting low-fidelity solution is then upscaled to a high-fidelity map via an enhanced autencoder.
@@ 5.2 Quality of high-fidelity solutions from studied architectures
Figure 15: Prediction results of thermal diffusion using the interpolation approach across three test scenarios.
@@ 6 Conclusion
Method Interpolation approach Feed Forward Neural Network MEA, Type 1 MEA, Type 2 Classical U-Net
@@
Possible disadvantages low accuracy for sharp solution low accuracy for sharp solution Network complexity Network complexity Network complexity / training cost
Main advantage Easy implementation Easy network High accuracy High accuracy Accuracy
@@ Acknowledgments
The authors are grateful to the Zentrum für Digitalisierungs- und Technologieforschung der Bundeswehr (dtec.bw) for their financial support.
@@ Appendix A Additional Results
Figure 28: Prediction results of heat flux using the MEA-Type1 model, featuring a twelve convolutional operation in encoder section.

### v2

@@ Abstract
In this study, we develop a novel multi-fidelity deep learning approach that transforms low-fidelity solution maps into high-fidelity ones by incorporating parametric space information into a standard autoencoder architecture. This method’s integration of parametric space information significantly reduces the need for training data to effectively predict high-fidelity solutions from low-fidelity ones. In this study, we examine a two-dimensional steady-state heat transfer analysis within a highly heterogeneous materials microstructure. The heat conductivity coefficients for two different materials are condensed from a $101\times 101$ grid to smaller grids. We then solve the boundary value problem on the coarsest grid using a pre-trained physics-informed neural operator network known as Finite Operator Learning (FOL). The resulting low-fidelity solution is subsequently upscaled back to a $101\times 101$ grid using a newly designed enhanced autoencoder. The novelty of the developed enhanced autoencoder lies in the concatenation of heat conductivity maps of different resolutions to the decoder segment in distinct steps. Hence the developed algorithm is named microstructure-embedded autoencoder (MEA). We compare the MEA outcomes with those from finite element methods, the standard U-Net, and various other upscaling techniques, including interpolation functions and feedforward neural networks (FFNN). Our analysis shows that MEA outperforms these methods in terms of computational efficiency and error on test cases. As a result, the MEA serves as a potential supplement to neural operator networks, effectively upscaling low-fidelity solutions to high fidelity while preserving critical details often lost in traditional upscaling methods, particularly at sharp interfaces like those seen with interpolation.
@@ 1 Introduction
To address the challenges associated with multi-fidelity approaches, we introduced a novel architecture called the Microstructure-Embedded Autoencoder (MEA). In this architecture, the low-fidelity solution map is first obtained by a solver on the reduced parametric space. Afterwards, the resulting low-fidelity solution is then upscaled to a high-fidelity map via an enhanced autoencoder.
@@ 5.2 Quality of high-fidelity solutions from studied architectures
Figure 15: Prediction results of thermal diffusion using the interpolation approach across three test scenarios. The contour plot in this and other similar figures shows the spatial distribution of the temperature.
@@ 6 Conclusion
Method Interpolation approach Feed Forward Neural Network MEA, Type 1    (this work) MEA, Type 2    (this work) Classical U-Net
@@
Possible disadvantages Poor accuracy for sharp solution Poor accuracy for sharp solution Network complexity Network complexity Network complexity / training cost
Main advantage Easy implementation Easy network architecture High accuracy even for sharp solution using moderate amount of data High accuracy even for sharp solution using moderate amount of data Accuracy (upon feeding enough data)
@@ Acknowledgments
The authors are grateful to the Zentrum für Digitalisierungs- und Technologieforschung der Bundeswehr (dtec.bw) for their financial support. The author Shahed Rezaei would like to thank the Deutsche Forschungsgemeinschaft (DFG) for the funding support provided to develop the present work in the project Cluster of Excellence “Internet of Production” (project: 390621612).
@@ Appendix A Additional Results
Figure 28: Prediction results of heat flux using the MEA-Type1 model, featuring a twelve convolutional operation in the encoder section.